OpenStudy (anonymous):

Does anyone understand imaginary numbers enough to explain them to me?? I really need to pass but I don't have a clue :(

terenzreignz (terenzreignz):

Well let's find out :) Do you have something specific that you're having a hard time understanding?

terenzreignz (terenzreignz):

Well, might as well cover the basics :D The root of all imaginary numbers is the constant i, defined as such... \[\Large i^2 = -1\]or \[\Large i = \sqrt{-1}\]

OpenStudy (anonymous):

Well for example (3+7x)(2i-19) and to factor it is really hard to understand

terenzreignz (terenzreignz):

that's x? not i?

OpenStudy (anonymous):

sorry meant i

OpenStudy (anonymous):

Really stressed out

terenzreignz (terenzreignz):

Okay. Don't forget that the constant i behaves like any normal variable. So you can FOIL it as you normally do :)

OpenStudy (anonymous):

Thank you :)

terenzreignz (terenzreignz):

That was it? Could you post your answer? :)

OpenStudy (anonymous):

Not exactly :/

OpenStudy (anonymous):

I understand it but yet don't if that makes any sense.

terenzreignz (terenzreignz):

FOIL method. could you multiply these binomials? \[\Large (7x+3)(2x-19)\]

terenzreignz (terenzreignz):

Remember:\[\Large (ax+b)(cx+d)= acx^2 + (bc+da)x + bd\]

OpenStudy (anonymous):

I can barely factor or do any algebraic thing...I was just passed to the next grade without any help and now im getting help with it and it is so confusing

terenzreignz (terenzreignz):

Well you need to practice :) What say we take an alternative way. Distributing. (3+7i)(2i-19) Let's distribute (3+7i) over (2i - 19) You get 2i(3+7i) - 19(3+7i) yes?

OpenStudy (anonymous):

Wait how exactly did you get the 2i and the 19 outside of the parenthese

terenzreignz (terenzreignz):

:) If it were a(2i - 19) You wouldn't hesitate to distribute the a over the (2i - 19) right? :) 2i(a) - 19(a)

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, you can treat the (3+7i) as the "a" and distribute it over the 2i - 19 all the same :)

OpenStudy (anonymous):

That sounds complicated

terenzreignz (terenzreignz):

but it isn't :) Just replace (a) with (3+7i)

OpenStudy (anonymous):

okay so if i were given an equation equal to zero and told to finnd the intercepts how would I if they involve i?

terenzreignz (terenzreignz):

example?

OpenStudy (anonymous):

Thank you for everything :)

OpenStudy (jhannybean):

O_o